### v1

@@ Abstract
Drug discovery typically consists of multiple steps, including identifying a target protein key to a disease’s etiology, validating that interacting with this target could prevent symptoms or cure the disease, discovering a small molecule or biologic therapeutic to interact with it, and optimizing the candidate molecule through a complex landscape of required properties. Drug discovery related tasks often involve prediction and generation while considering multiple entities that potentially interact, which poses a challenge for typical AI models. For this purpose we present MAMMAL - Molecular Aligned Multi-Modal Architecture and Language - a method that we applied to create a versatile multi-task foundation model ibm/biomed.omics.bl.sm.ma-ted-458m that learns from large-scale biological datasets (2 billion samples) across diverse modalities, including proteins, small molecules, and genes. We introduce a prompt syntax that supports a wide range of classification, regression, and generation tasks. It allows combining different modalities and entity types as inputs and/or outputs. Our model handles combinations of tokens and scalars and enables the generation of small molecules and proteins, property prediction, and transcriptomic lab test predictions. We evaluated the model on 11 diverse downstream tasks spanning different steps within a typical drug discovery pipeline, where it reaches new SOTA in 9 tasks and is comparable to SOTA in 2 tasks. This performance is achieved while using a unified architecture serving all tasks, in contrast to the original SOTA performance achieved using tailored architectures.
@@
Figure 1: (A) We introduce a multi-aligned model pretrained on six datasets, each containing tens to hundreds of millions of data points. These data points include protein sequences, small molecules, and gene expression profiles, with a combined sample size of 2 billion. (B) The multi-aligned model combines flexible encoder-only and encoder-decoder components. It takes sequences as input, which may contain any combination of tokens and scalar elements, processed by an encoder stack consisting of self-attention blocks. In encoder-only mode, a dedicated token prediction head outputs logits for token predictions, with an optional scalar prediction head for scalar outputs. In encoder-decoder mode, residual connections inject features from the encoder’s final hidden layer into each decoder layer, and a decoder-specific prediction head outputs the final logits. (C) Diverse downstream tasks performed by the multi-aligned model, mapped to their contributions within the steps of a typical drug discovery pipeline. (D) Diverse downstream tasks performed by the multi-aligned model, categorized by data type used in the fine-tuning process. (E) Performance of the multi-aligned model across a diverse set of tasks compared to SOTA.
@@ 1 Introduction
In this work, we introduce the MAMMAL (Molecular Aligned Multi-Modal Architecture and Language) method, and develop a multi-aligned foundation model paired with a prompt syntax that integrates multiple data domains to support a wide range of drug discovery tasks. The multi-aligned model, which enables aligning multiple entities into a single prompt, has been extensively pre-trained on 2 billion samples from diverse datasets, using auxiliary tasks of mask infilling, denoising, generation, and classification. The multi-aligned model is compatible with both encoder-decoder and encoder-only architectures and effectively incorporates numerical values through continuous token embedding, enhancing numerical precision and reducing vocabulary size. We rigorously evaluate the multi-aligned model across 11 downstream tasks - spanning classification, regression, and generation - covering key stages of the drug discovery pipeline across three primary domains: small molecules, proteins, and gene expression profiles. The multi-aligned model achieves state-of-the-art performance in 9 tasks and matches top performance in the remaining 2 tasks. The model is publicly available on https://huggingface.co/ibm/biomed.omics.bl.sm.ma-ted-458m under the name ibm/biomed.omics.bl.sm.ma-ted-458m
@@ 3.1 Cell Type Annotation
Cell type prediction enables researchers to distinguish between different cell populations, such as those associated with various diseases [9, 10, 11, 12]. It is also crucial for understanding how diseases or drugs affect different cell types. In recent years, a variety of methods have been developed for this task, including approaches based on marker genes, correlation-based techniques, and annotation using classification [37]. Recent advances in transformer-based and large-scale foundation models [38, 28, 39] have improved performance by utilizing the full list of genes as input, in contrast to earlier methods where gene selection was based on highly Variable genes (HVG).
The input for this task commonly consists of gene expression (GE) values from single-cell RNA-seq data. The benchmark we used is based on the Zheng68k dataset [40], which is derived from human peripheral blood mononuclear cells (PBMCs) and is widely used for evaluating cell-type annotation performance. The dataset contains 68,579 cells across 11 cell types and originally included 32,738 genes. Preprocessing involved normalization, log-transformation of expression values and followed filtering out non-expressed genes, leaving around 20,387 genes. Similar to the approach in [41], our model uses a ranked list of expressed gene names, ordered by their expression levels, as input. The label to predict is provided in the cell ontology format “CL:NNNNNN” (see Table LABEL:table_task_queries).
@@ 3.8 Drug-Target Interaction
Predicting drug-target binding affinity plays a crucial role in the early stages of drug discovery. Traditionally, binding affinities are measured through high-throughput screening experiments, which, while accurate, are resource-intensive and limited in their scalability to evaluate large sets of drug candidates. In this task, we focus on predicting binding affinities using pKd, the negative logarithm of the dissociation constant, which reflects the strength of the interaction between a small molecule (drug) and a protein (target). We utilize the PEER(Protein sEquence undERstanding) benchmark [36] for drug-target interaction (DTI) prediction. This benchmark leverages data from the BindingDB dataset [68], with a specific test split that holds out four protein classes - estrogen receptor (ER), G-protein-coupled receptors (GPCR), ion channels, and receptor tyrosine kinases - for assessing generalization performance on unseen classes.
For model fine-tuning, we conducted hyperparameter optimization, selecting an initial learning rate of 0.0004, with no dropout and no weight decay. We standardized the pKd values based on the mean and standard deviation of the training set. For evaluation, we transformed the predicted values back to their original scale. As shown in Table 2, our model achieved an average NRMSE of 0.906, demonstrating a solid improvement of 3.8% over the SOTA reported by [36].
@@ 4 Discussion
Drug discovery is often hindered by a lack of large, high-quality datasets, particularly in biomedicine, where data generation and collection are costly and face challenges related to privacy, lack of standardization, and ethical constraints. A planned extension of our approach involves the support of free-text input, enabling pretraining on extensive biomedical text sources like PubMed as well as the incorporation of free-text segments into prompts of fine-tuned models. Free text provides rich, nuanced context and allows users greater flexibility beyond predefined formats. The integration of large language models, which capture human narratives, with our newly introduced multi-aligned model-demonstrated to effectively encode complex interactions within protein complexes, between proteins and small molecules, and between small molecules and cells, including their gene expression levels-holds promise of creating a model capable of uncovering new scientific knowledge.
@@ 5 Data Availability
PPI. The Weber TCR binding dataset was downloaded from [62], and the HER2 antibody-antigen binding dataset was taken from the original paper [61] github repository [70].
Pretraining. ibm/biomed.omics.bl.sm.ma-ted-458m was pre-trained over OAS [23], Uniref90 [71], Zinc [24], PubChem [25] and CELLxGENE [26]. Appendix C describes the pre-processing steps applied.
@@ Appendix A Architecture - additional details
A user prompt, usually expressed as a single text line, is processed into 2 input sequences: a. Input token IDs, which is a sequence of integer values representing tokens in the vocabulary b. a sequence of inputs scalars (by convention, containing NaNs for positions for which no input scalar is provided).

### v2

@@ Abstract
Drug discovery typically consists of multiple steps, including identifying a target protein key to a disease’s etiology, validating that interacting with this target could prevent symptoms or cure the disease, discovering a small molecule or biologic therapeutic to interact with it, and optimizing the candidate molecule through a complex landscape of required properties. Drug discovery related tasks often involve prediction and generation while considering multiple entities that potentially interact, which poses a challenge for typical AI models. For this purpose we present MAMMAL - Molecular Aligned Multi-Modal Architecture and Language - a method that we applied to create a versatile multi-task multi-align foundation model that learns from large-scale biological datasets (2 billion samples) across diverse modalities, including proteins, small molecules, and genes. We introduce a prompt syntax that supports a wide range of classification, regression, and generation tasks. It allows combining different modalities and entity types as inputs and/or outputs. Our model handles combinations of tokens and scalars and enables the generation of small molecules and proteins, property prediction, and transcriptomic lab test predictions. We evaluated the model on 11 diverse downstream tasks spanning different steps within a typical drug discovery pipeline, where it reaches new SOTA in 9 tasks and is comparable to SOTA in 2 tasks. This performance is achieved while using a unified architecture serving all tasks, in contrast to the original SOTA performance achieved using tailored architectures.
@@
Figure 1: (A) We introduce a multi-align model pretrained on six datasets, each containing tens to hundreds of millions of data points. These data points include protein sequences, small molecules, and gene expression profiles, with a combined sample size of 2 billion. (B) The multi-align model combines flexible encoder-only and encoder-decoder components. It takes sequences as input, which may contain any combination of tokens and scalar elements, processed by an encoder stack consisting of self-attention blocks. In encoder-only mode, a dedicated token prediction head outputs logits for token predictions, with an optional scalar prediction head for scalar outputs. In encoder-decoder mode, residual connections inject features from the encoder’s final hidden layer into each decoder layer, and a decoder-specific prediction head outputs the final logits. (C) Diverse downstream tasks performed by the multi-align model, mapped to their contributions within the steps of a typical drug discovery pipeline. (D) Diverse downstream tasks performed by the multi-align model, categorized by data type used in the fine-tuning process. (E) Performance of the multi-align model across a diverse set of tasks compared to SOTA.
@@ 1 Introduction
In this work, we introduce the MAMMAL (Molecular Aligned Multi-Modal Architecture and Language) method, and develop a multi-align foundation model paired with a prompt syntax that integrates multiple data domains to support a wide range of drug discovery tasks. The multi-align model, which enables aligning multiple entities into a single prompt, has been extensively pre-trained on 2 billion samples from diverse datasets, using auxiliary tasks of mask infilling, denoising, generation, and classification. The multi-align model is compatible with both encoder-decoder and encoder-only architectures and effectively incorporates numerical values through continuous token embedding, enhancing numerical precision and reducing vocabulary size. We rigorously evaluate the multi-align model across 11 downstream tasks - spanning classification, regression, and generation - covering key stages of the drug discovery pipeline across three primary domains: small molecules, proteins, and gene expression profiles. The multi-align model achieves state-of-the-art performance in 9 tasks and matches top performance in the remaining 2 tasks. The model is publicly available on https://huggingface.co/ibm/biomed.omics.bl.sm.ma-ted-458m under the name ibm/biomed.omics.bl.sm.ma-ted-458m
@@ 3.1 Cell Type Annotation
Cell type prediction enables researchers to distinguish between different cell populations, such as those associated with various diseases [9, 10, 11, 12]. It is also crucial for understanding how diseases or drugs affect different cell types. In recent years, a variety of methods have been developed for this task, including approaches based on marker genes, correlation-based techniques, and annotation using classification [37]. Recent advances in transformer-based and large-scale foundation models [38, 28, 39] have improved performance by utilizing the full list of genes as input, in contrast to earlier methods where gene selection was based on highly variable genes.
The input for this task commonly consists of gene expression (GE) values from single-cell RNA-seq data. The benchmark we used is based on the Zheng68k dataset [40], which is derived from human peripheral blood mononuclear cells and is widely used for evaluating cell-type annotation performance. The dataset contains 68,579 cells across 11 cell types and originally included 32,738 genes. Preprocessing involved normalization, log-transformation of expression values and followed filtering out non-expressed genes, leaving around 20,387 genes. Similar to the approach in [41], our model uses a ranked list of expressed gene names, ordered by their expression levels, as input. The label to predict is provided in the cell ontology format “CL:NNNNNN” (see Table LABEL:table_task_queries).
@@ 3.8 Drug-Target Interaction
Predicting drug-target binding affinity plays a crucial role in the early stages of drug discovery. Traditionally, binding affinities are measured through high-throughput screening experiments, which, while accurate, are resource-intensive and limited in their scalability to evaluate large sets of drug candidates. In this task, we focus on predicting binding affinities using $\textnormal{p}K_{D}$, the negative logarithm of the dissociation constant, which reflects the strength of the interaction between a small molecule (drug) and a protein (target). We utilize the PEER (Protein sEquence undERstanding) benchmark [36] for DTI prediction. This benchmark leverages data from the BindingDB dataset [68], with a specific test split that holds out four protein classes - estrogen receptor, G-protein-coupled receptors, ion channels, and receptor tyrosine kinases - for assessing generalization performance on unseen classes.
For model fine-tuning, we conducted hyperparameter optimization, selecting an initial learning rate of 0.0004, with no dropout and no weight decay. We standardized the $\textnormal{p}K_{D}$ values based on the mean and standard deviation of the training set. For evaluation, we transformed the predicted values back to their original scale. As shown in Table 2, our model achieved an average NRMSE of 0.906, demonstrating a solid improvement of 3.8% over the SOTA reported by [36].
@@ 4 Discussion
Drug discovery is often hindered by a lack of large, high-quality datasets, particularly in biomedicine, where data generation and collection are costly and face challenges related to privacy, lack of standardization, and ethical constraints. A planned extension of our approach involves the support of free-text input, enabling pretraining on extensive biomedical text sources like PubMed as well as the incorporation of free-text segments into prompts of fine-tuned models. Free text provides rich, nuanced context and allows users greater flexibility beyond predefined formats. The integration of large language models, which capture human narratives, with our newly introduced multi-align model-demonstrated to effectively encode complex interactions within protein complexes, between proteins and small molecules, and between small molecules and cells, including their gene expression levels-holds promise of creating a model capable of uncovering new scientific knowledge.
@@ 5 Data Availability
TCR Bind The Weber TCR binding dataset was downloaded from [62], and the HER2 antibody-antigen binding dataset was taken from the original paper [61] github repository [70].
Pretraining. ibm/biomed.omics.bl.sm.ma-ted-458m was pre-trained over OAS [23], UniProt [71], Zinc [24], PubChem [25], STRING [27] and CELLxGENE [26]. Appendix C describes the pre-processing steps applied.
@@ Appendix A Architecture - additional details
A user prompt, usually expressed as a single text line, is processed into two input sequences: a. Input token IDs, which is a sequence of integer values representing tokens in the vocabulary, and b. a sequence of inputs scalars (by convention, containing NaNs for positions for which no input scalar is provided).